### v1

# WeightedPose: Generalizable Cross-Pose Estimation via Weighted SVD

Abstract
We present a novel method for robotic manipulation tasks in human environments that require reasoning about the 3D geometric relationship between a pair of objects. Traditional end-to-end trained policies, which map from pixel observations to low-level robot actions, struggle to reason about complex pose relationships and have difficulty generalizing to unseen object configurations. To address these challenges, we propose a method that learns to reason about the 3D geometric relationship between objects, focusing on the relationship between key parts on one object with respect to key parts on another object. Our standalone model utilizes Weighted SVD to reason about both pose relationships between articulated parts and between free-floating objects. This approach allows the robot to understand the relationship between the oven door and the oven body, as well as the relationship between the lasagna plate and the oven, for example. By considering the 3D geometric relationship between objects, our method enables robots to perform complex manipulation tasks that reason about object-centric representations. We open source the code and demonstrate the results here  [^1] https://github.com/harryzhangOG/weighted-pose/tree/main.
1 Introduction
Many manipulation tasks require a robot to move an object to a location relative to another object. For example, a cooking robot may need to place a lasagna in an oven, place a pot on a stove, place a plate in a microwave, place a mug onto a mug rack, or place a cup onto a shelf. Understanding and placing objects in task-specific locations is a key skill for robots operating in human environments. Further, this skill should generalize to novel objects within the training categories, such as placing new trays into the oven or new mugs onto a mug rack.
A common approach in robot learning is to train a policy “end-to-end," mapping from pixel observations to low-level robot actions. However, end-to-end trained policies cannot easily reason about complex pose relationships such as the ones described above, and they have difficulty generalizing to unseen object configurations. In contrast, we propose a method that learns to reason about the 3D geometric relationship between a pair of objects. For the type of tasks defined above, the robot needs to reason about the relationship between key parts on one object with respect to key parts on another object. Specifically, the robot needs to reason about both the relationship between the oven door and the oven body, as well as the relationship between the lasagna plate and the oven. We propose a standalone model that utilizes Weighted SVD to reason about both pose relationship between articulated parts and between free-floating objects.
2 Background
Object Pose Estimation: Pose estimation is the task of detecting and inferring the 6DoF pose of an object, which includes its position and orientation, with respect to some previously defined object reference frame (19; 28; 36; 10; 11; 32). Recent work (20; 27; 33; 21) proposed to use 3D semantic keypoints as an alternative form of object representation. While keypoint-based methods can generalize within an object class, they require a significant amount of hand annotated data or access to a simulated version of the task to learn to estimate the keypoint locations. In contrast, our method is able to learn from just 10 real-world demonstrations. Another approach is to use dense embeddings, such as Dense Object Nets (DON) (9) and Neural Descriptor Fields (NDF) (31; 7), which achieve generalization across classes by predicting dense embeddings in the observation and matching them to embeddings of the demonstration objects. However, DON (9) and NDF (31) assume that the target object is moved relative to a static reference object in a “known canonical configuration" (e.g. the pose of the mug rack in NDF (31) is assumed to be known and fixed). In contrast, our method reasons about the geometric relationship between a pair of objects and hence does not need to assume a static environment.
Articulated Object Manipulation: Manipulation of articulated objects and other objects with non-rigid properties remains an open research area due to the objects’ complex geometries and kinematics. Previous work proposed manipulating such objects by hand-designed analytical methods, such as the immobilization of a chain of hinged objects by (5; 29; 1; 2; 39). Berenson et al. (3) proposed a planning framework for manipulation under kinematic constraints. Katz et al. (15) proposed a method to learn such manipulation policies in the real-world using a grounded relational representation learned through interaction. With the development of larger-scale datasets of articulated objects such as the PartNet dataset by Mo et al. (22) and Partnet-Mobility by Xiang et al. (35); Sim et al. (30); Jin et al. (14), several works have proposed learning methods based on large-scale simulation and supervised visual learning. Mo et al. (23) proposed to learn articulation manipulation policies through large-scale simulation and visual affordance learning. Xu et al. (37) proposed a system that learns articulation affordances as well as an action scoring module, which can be used to articulate objects. Several works have focused specifically on visual recognition and estimation of articulation parameters, learning to predict the pose (42; 45; 43; 38; 34; 12; 16) and identify articulation parameters (13; 41; 44) to obtain action trajectories. Moreover, (24; 4; 6; 17; 18; 8) tackle the problem using statistical motion planning.
3 Method
Figure 1: Unified Weighted Pose architecture. The model first takes as input a point cloud, and then learns to predict a weight for the point cloud. This weight is used in the downstream SVD module to combine the GoalFlow and TAX-Pose outputs.
We first slightly modify the FlowBot 3D training objective. Instead of defining an instantaneous motion vector field as the training data directly, we make the network learn to output a flow field to the completely open state directly. Since this version of FlowBot 3D learns to output a dense representation to the goal state directly, we call this model Goal Flow.
Formally, given a point cloud $\{\mathbf{p}_{i}\}\;\forall i\in\{1,\dots,N\}$ this Goal Flow model outputs a dense flow field $F\in\mathbb{R}^{N\times 3}$, where each flow vector $\delta_{i}\in\mathbb{R}^{3}$ in the $F$ represents a goal flow vector such that point $p_{i}+\delta_{i}$ is in the fully open goal state. Ideally, this model should be deployed exclusively for articulated objects in that Goal Flow was shown to achieve suboptimal performances in (25; 26).
To combine the Goal Flow model with TAX-Pose, we also make the Goal Flow network output an auxiliary weight $w\in\mathbb{R}$, which assigns weight $w$ to Goal Flow and $1-w$ to TAX-Pose.
For the TAX-Pose component of the Weighted Pose unified architecture, we do not make any significant modifications and ideally the TAX-Pose model should be deployed for free-floating objects exclusively.
In TAX-Pose, we have:
$\mathcal{J}(\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}})=\sum_{i=1}^{N_{{\mathcal{A}}}}\alpha_{i}^{{\mathcal{A}}}||\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}~{}\mathbf{p}_{i}^{{\mathcal{A}}}-\tilde{\mathbf{v}}_{i}^{{\mathcal{A}}}||_{2}^{2}+\sum_{i=1}^{N_{{\mathcal{B}}}}\alpha_{i}^{{\mathcal{B}}}||\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}^{-1}~{}\mathbf{p}_{i}^{{\mathcal{B}}}-\tilde{\mathbf{v}}_{i}^{{\mathcal{B}}}||_{2}^{2},$
where
$\mathbf{A}=\begin{bmatrix}\mathbf{P}_{{\mathcal{A}}}^{*\top}&\mathbf{\tilde{V}}_{{\mathcal{B}}}^{*\top}\end{bmatrix},\;\;\mathbf{B}=\begin{bmatrix}\mathbf{\tilde{V}}_{{\mathcal{A}}}^{*\top}&\mathbf{P}_{{\mathcal{B}}}^{*\top}\end{bmatrix}^{\top},\;\;\bm{\Gamma}=\text{diag}\left(\begin{bmatrix}\bm{\alpha}_{{\mathcal{A}}}&\bm{\alpha}_{{\mathcal{B}}}\end{bmatrix}\right)$
Now we want to add another term in the SVD step. Specifically, we want the network to learn which one of the two models, TAX-Pose and Goal-Flow, is more important based on point cloud observations. Thus, we want a SVD step that incorporates the TAX-Pose residual, weighted by $(1-w)$ and Goal-Flow weighted by $w$. So the new SVD formulation becomes:
$\mathcal{J}(\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}})$ $=(1-w)\left[\sum_{i=1}^{N_{{\mathcal{A}}}}\alpha_{i}^{{\mathcal{A}}}||\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}~{}\mathbf{p}_{i}^{{\mathcal{A}}}-\tilde{\mathbf{v}}_{i}^{{\mathcal{A}}}||_{2}^{2}+\sum_{i=1}^{N_{{\mathcal{B}}}}\alpha_{i}^{{\mathcal{B}}}||\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}^{-1}~{}\mathbf{p}_{i}^{{\mathcal{B}}}-\tilde{\mathbf{v}}_{i}^{{\mathcal{B}}}||_{2}^{2}\right]$
$+w\sum_{i=1}^{N_{{\mathcal{A}}}}||\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}~{}\mathbf{p}_{i}^{{\mathcal{A}}}-(\mathbf{p}_{i}^{{\mathcal{A}}}+\delta^{\mathcal{A}}_{i})||_{2}^{2},$
where $\delta^{\mathcal{A}}_{i}$ is the $i-$th point’s goal flow. To make this approach viable we optimize $R$ and $t$ in $\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}$ separately:
$J$ $=\sum_{i}^{N_{\mathcal{A}}}\left((1-w)\alpha_{i}\|R\mathbf{p}_{i}^{\mathcal{A}}+t-\mathbf{\tilde{v}}_{i}^{\mathcal{A}}\|^{2}+w\|R\mathbf{p}_{i}^{\mathcal{A}}+t-\mathbf{p}_{i}^{\mathcal{A}}+\delta_{i}^{\mathcal{A}}\|^{2}\right)$
$+\sum_{i}^{N_{\mathcal{B}}}(1-w)\alpha_{i}\|R^{-1}\mathbf{p}_{i}^{\mathcal{B}}-t-\mathbf{\tilde{v}}_{i}^{\mathcal{B}}\|^{2}$
We first solve for the optimal translation $t^{*}$:
$\frac{\partial J}{\partial t}$ $=0$
$=\sum_{i}^{N_{\mathcal{A}}}\left(2(1-w)\alpha_{i}\left(t+R\mathbf{p}_{i}^{\mathcal{A}}-\mathbf{\tilde{v}}_{i}^{\mathcal{A}}\right)+2w\left(t+R\mathbf{p}_{i}^{\mathcal{A}}-\mathbf{p}_{i}^{\mathcal{A}}-\delta_{i}^{\mathcal{A}}\right)\right)$
$+2(1-w)\sum_{i}^{N_{\mathcal{B}}}\alpha_{i}\left(R^{-1}\mathbf{p}_{i}^{\mathcal{B}}-t-\mathbf{\tilde{v}}_{i}^{\mathcal{B}}\right)$
$=\sum_{i}^{N_{\mathcal{A}}}\left((2-2w)\alpha_{i}+2w\right)t+\left((2-2w)\alpha_{i}+2w\right)R\mathbf{p}_{i}^{\mathcal{A}}-2w\delta_{i}^{\mathcal{A}}-(2-2w)\alpha_{i}\mathbf{\tilde{v}}_{i}^{\mathcal{A}}-2w\mathbf{p}_{i}^{\mathcal{A}}$
$+\sum_{i}^{N_{\mathcal{B}}}-(2-2w)\alpha_{i}t+(2-2w)\alpha_{i}(R^{-1}\mathbf{p}_{i}^{\mathcal{B}}-\mathbf{\tilde{v}}_{i}^{\mathcal{B}})$
$=\left[\sum_{i}^{N_{\mathcal{A}}}[(2-2w)\alpha_{i}+2w]+\sum_{i}^{N_{\mathcal{B}}}-(2-2w)\alpha_{i}\right]t$
$+\sum_{i}^{N_{\mathcal{A}}}\left((2-2w)\alpha_{i}+2w\right)R\mathbf{p}_{i}^{\mathcal{A}}-2w\delta_{i}^{\mathcal{A}}-(2-2w)\alpha_{i}\mathbf{\tilde{v}}_{i}^{\mathcal{A}}-2w\mathbf{p}_{i}^{\mathcal{A}}$
$+\sum_{i}^{N_{\mathcal{B}}}(2-2w)\alpha_{i}(R^{-1}\mathbf{p}_{i}^{\mathcal{B}}-\mathbf{\tilde{v}}_{i}^{\mathcal{B}})$
$t^{*}$ $=-\bigg{(}\sum_{i}^{N_{\mathcal{A}}}\left((2-2w)\alpha_{i}+2w\right)R\mathbf{p}_{i}^{\mathcal{A}}-2w\delta_{i}^{\mathcal{A}}-(2-2w)\alpha_{i}\mathbf{\tilde{v}}_{i}^{\mathcal{A}}-2w\mathbf{p}_{i}^{\mathcal{A}}$
$+\sum_{i}^{N_{\mathcal{B}}}(2-2w)\alpha_{i}(R^{-1}\mathbf{p}_{i}^{\mathcal{B}}-\mathbf{\tilde{v}}_{i}^{\mathcal{B}})\bigg{)}\bigg{/}\left[\sum_{i}^{N_{\mathcal{A}}}[(2-2w)\alpha_{i}+2w]+\sum_{i}^{N_{\mathcal{B}}}-(2-2w)\alpha_{i}\right]$
Further simplifying, we have:
$t^{*}$ $=\frac{\color[rgb]{1,0,0}(1-w)\sum_{i}^{N_{\mathcal{A}}}\alpha_{i}^{\mathcal{A}}(\mathbf{\tilde{v}}_{i}^{\mathcal{A}}-R\mathbf{p}_{i}^{\mathcal{A}})}{\color[rgb]{1,0,0}(1-w)\sum_{i}^{N_{\mathcal{A}}}\alpha_{i}^{\mathcal{A}}+\color[rgb]{0,0,1}w\sum_{i}^{N_{\mathcal{A}}}1+\color[rgb]{.75,0,.25}(1-w)\sum_{i}^{N_{\mathcal{B}}}\alpha_{i}^{\mathcal{B}}}$
$+\frac{\color[rgb]{0,0,1}w\sum_{i}^{N_{\mathcal{A}}}(\mathbf{p}_{i}^{\mathcal{A}}+\delta_{i}^{\mathcal{A}})-R\mathbf{p}_{i}^{\mathcal{A}}}{\color[rgb]{1,0,0}(1-w)\sum_{i}^{N_{\mathcal{A}}}\alpha_{i}^{\mathcal{A}}+\color[rgb]{0,0,1}w\sum_{i}^{N_{\mathcal{A}}}1+\color[rgb]{.75,0,.25}(1-w)\sum_{i}^{N_{\mathcal{B}}}\alpha_{i}^{\mathcal{B}}}$
$+\frac{\color[rgb]{.75,0,.25}(1-w)\sum_{i}^{N_{\mathcal{B}}}\alpha_{i}^{\mathcal{B}}(\mathbf{\tilde{v}}_{i}^{\mathcal{B}}-R^{-1}\mathbf{p}_{i}^{\mathcal{B}})}{\color[rgb]{1,0,0}(1-w)\sum_{i}^{N_{\mathcal{A}}}\alpha_{i}^{\mathcal{A}}+\color[rgb]{0,0,1}w\sum_{i}^{N_{\mathcal{A}}}1+\color[rgb]{.75,0,.25}(1-w)\sum_{i}^{N_{\mathcal{B}}}\alpha_{i}^{\mathcal{B}}}$
Note here that we colorcode the expression here. Intuitively, the resulting translation is a weighted sum of three translation terms. The red color represents the action object’s translation via TAX-Pose, the blue color represents the action object’s translation via GoalFlow, and the purple color represents the anchor object’s translation via TAX-Pose.
We can then construct the matrices as follows:
$\mathbf{A}=\begin{bmatrix}\mathbf{P}_{{\mathcal{A}}}^{\top}&\mathbf{\tilde{V}}_{{\mathcal{B}}}^{\top}&\mathbf{P}_{{\mathcal{A}}}^{\top}\end{bmatrix},\;\;\mathbf{B}=\begin{bmatrix}\mathbf{\tilde{V}}_{{\mathcal{A}}}^{\top}&\mathbf{P}_{{\mathcal{B}}}^{\top}&\mathbf{P}_{{\mathcal{A}}}^{\top}+\Delta_{\mathcal{A}}\end{bmatrix}^{\top}$
$\bm{\Gamma}=\begin{bmatrix}(1-w)\cdot\bm{\alpha}_{{\mathcal{A}}}&0&0\\ 0&(1-w)\cdot\bm{\alpha}_{{\mathcal{B}}}&0\\ 0&0&w\end{bmatrix}$
where $\Delta_{\mathcal{A}}$ is the de-meaned goal flow field. Note here that everything here is NOT de-meaned.
We then solve for the SVD:
$\bm{U\Sigma V^{\top}}=\text{svd}(\bm{A\Gamma B^{\top}})$
and then we solve for rotation matrix $R$. Plugging it back into the $t^{*}$ equations we get $t$.
To train the WeightedPose, we use a set of losses defined below, which are similar to those in (25). We assume we have access to a set of demonstrations of the task, in which the action and anchor objects are in the target relative pose such that $\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}=\mathbf{I}$.
Point Displacement Loss: Instead of directly supervising the rotation and translation (as is done in DCP), we supervise the predicted transformation using its effect on the points. For this loss, we take the point clouds of the objects in the demonstration configuration, and transform each cloud by a random transform, $\mathbf{\hat{P}}_{\mathcal{A}}=\mathbf{T}_{\alpha}\mathbf{P}_{\mathcal{A}}$, and $\mathbf{\hat{P}}_{\mathcal{B}}=\mathbf{T}_{\beta}\mathbf{P}_{\mathcal{B}}$. This would give us a ground truth transform of $\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}^{GT}=\mathbf{T}_{\beta}\mathbf{T}_{\alpha}^{-1}$; the inverse of this transform would move object $\mathcal{B}$ to the correct position relative to object $\mathcal{A}$. Using this ground truth transform, we compute the MSE loss between the correctly transformed points and the points transformed using our prediction.
$\mathcal{L}_{\mathrm{disp}}=\left\|\mathbf{T}_{\mathcal{A}\mathcal{B}}\mathbf{P}_{\mathcal{A}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}^{GT}\mathbf{P}_{\mathcal{A}}\right\|^{2}+\left\|\mathbf{T}_{\mathcal{A}\mathcal{B}}^{-1}\mathbf{P}_{\mathcal{B}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}^{GT-1}\mathbf{P}_{\mathcal{B}}\right\|^{2}$ (1)
Direct Correspondence Loss. While the Point Displacement Loss best describes errors seen at inference time, it can lead to correspondences that are inaccurate but whose errors average to the correct pose. To improve these errors we directly supervise the learned correspondences $\tilde{V}_{\mathcal{A}}$ and $\tilde{V}_{\mathcal{B}}$:
$\mathcal{L}_{\mathrm{corr}}=\left\|\mathbf{\tilde{V}}_{{\mathcal{A}}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}^{GT}\mathbf{P}_{\mathcal{A}}\right\|^{2}+\left\|\mathbf{\tilde{V}}_{{\mathcal{B}}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}^{GT-1}\mathbf{P}_{\mathcal{B}}\right\|^{2}.$ (2)
Correspondence Consistency Loss. Furthermore, a consistency loss can be used. This loss penalizes correspondences that deviate from the final predicted transform. A benefit of this loss is that it can help the network learn to respect the rigidity of the object, while it is still learning to accurately place the object. Note, that this is similar to the Direct Correspondence Loss, but uses the predicted transform as opposed to the ground truth one. As such, this loss requires no ground truth:
$\mathcal{L}_{\mathrm{cons}}=\left\|\mathbf{\tilde{V}}_{{\mathcal{A}}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}\mathbf{P}_{\mathcal{A}}\right\|^{2}+\left\|\mathbf{\tilde{V}}_{{\mathcal{B}}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}^{-1}\mathbf{P}_{\mathcal{B}}\right\|^{2}.$ (3)
Direct SE(3) Transformation Loss. We also define a loss directly for the transformation outputted in $SE(3)$, supervising between the SVD output pose and ground-truth pose.
$\mathcal{L}_{\text{tf}}=||\mathbf{T}_{\mathcal{A}\mathcal{B}}-\mathbf{T}_{\mathcal{A}\mathcal{B}}^{GT}||_{F}$ (4)
4 Experiments
We define a PartNet-Mobility Placement task as placing a given action object relative to an anchor object based on a semantic goal position. This would require the model to be able to output cross-pose for both the articulated parts and the free-floating objects. We select a set of household furniture objects from the PartNet-Mobility dataset (35) and a set of small rigid objects released with the Ravens simulation environment (40). When the model is required to estimate goal pose for the articulated part, following the terminology in (25), the action object would be the articulated part (e.g. doors, drawers), and the anchor object would be the static body of the furniture. In contrast, when the model is estimating goal pose for free floating objects, then it becomes the same setting as in (25).
Figure 2: Task illustration. The model needs to first output goal pose for opening the oven door, and then output goal pose for putting the block inside the oven.
We compare the proposed method to several baselines. First, we compare against TAX-Pose trained on both free-floating and articulated objects as well as Goal-Flow trained on both categories of objects. We then compare against an oracle model that classifies if the objects of interest are free-floating or articulated, and deploys the model accordingly.
We describe the baselines in details:
  • Goal Flow Pretrained (GF Pretrained): Pretrain Goal Flow on articulated objects only and test on both articulated and free-floating objects.
  • TAX-Pose Pretrained (TP Pretrained): Pretrain TAX-Pose on free-floating objects only and test on both articulated and free-floating objects.
  • Weighted Pose Original Loss (WP OG Loss): Weighted Pose trained and test on both free-floating and articulated objects using the original TAX-Pose loss.
  • Weighted Pose Post-SVD Loss (WP Post-SVD): Weighted Pose trained and test on both free-floating and articulated objects but using the post-SVD TAX-Pose loss.
  • Weighted Pose Post-SVD & Transformation Loss (WP Post-SVD): Weighted Pose trained and test on both free-floating and articulated objects but using both the post-SVD TAX-Pose loss and direct transformation loss. Where the transformation loss is the MSE between the predicted $SE(3)$ transformation and the ground-truth $SE(3)$ transformation.
We use the same PartNet-Mobility dataset as in TAX-Pose to evaluate the system. For semantic tasks, we will only select the “In” task as it is the most intuitive task. However, other than just performing evaluation on the free-floating objects, we will also evaluate the method on opening the articulate objects. Under this formulation, the action object is the articulated part and the anchor object is the static body. We will measure rotational error. translational error, and per-point MSE in our experiments. We follow the same train-val split as in TAX-Pose’s PartNet-Mobility dataset.
In Table 1, Goal Flow Pretrained baseline fails on free-floating objects and excels on articulated objects. Similarly for TAX-Pose pretrained, since it was trained on free-floating objects, it does do well on articulated objects, as the high rotational and translational errors indicate.
We next evaluate the variations of WeightedPose, which differ in the training paradigms. Using Weighted Pose Original Loss, which is trained using the original TAX-Pose loss, we are able to achieve better performance on both articulated and free-floating objects. Interestingly, the results we achieve using Weighted Pose Original Loss for articulated objects are better than the results from using Goal Flow pretrained on articulated objects.
While ideally, the $w$ learned in this method should effectively act as a classifier of the input object (1 for articulated objects, 0 for free-floating objects). Intuitively, given a perfect $w$, the performance of WeightedPose would be upper-bounded by Goal Flow’s performance on articulated objects and by TAX-Pose on free-floating objects. However, a weighted combination of the two results due to the imperfect learned $w$ weight could potentially correct the mistake made by each model by summing the results with a weighted sum. This may explain why WeightedPose sometimes performs better than its hypothesized upper bound on some objects.
Lastly, results suggest that the original loss used in TAX-Pose yields the best overall performance. However, it is worth noting that by using post-SVD loss during training, we are able to achieve lower translational error in test time. Interestingly, by introducing a direct $SE(3)$ supervision, the results degrade marginally.
GF Pretrained TP Pretrained WP OG Loss WP Post-SVD WP Post SVD + T
Metrics FF Art FF Art FF Art FF Art FF Art
Train Rot err 31.61 5.16 2.61 55.78 11.69 3.15 13.14 3.28 14.04 5.52
Trans err 1.21 0.16 0.04 0.98 0.14 0.08 0.21 0.07 0.21 0.07
PP MSE 1.04 0.04 0.01 0.85 0.07 0.05 0.09 0.04 0.11 0.08
Val Rot err 35.5 9.14 9.87 59.73 11.22 9.01 14.13 9.71 13.06 10.03
Trans err 1.3 0.19 0.18 0.99 0.26 0.15 0.22 0.18 0.23 0.18
PP MSE 1.07 0.1 0.15 0.82 0.16 0.11 0.17 0.12 0.18 0.11
Table 1: Weighted Pose Results: We compare Rotation Error, Translational Error, and Per-Point MSE for both training and validation objects.
5 Conclusions and Future Work
In conclusion, we have presented a method to combine the two architectures using weighted SVD. While this model is more of a proof-of-concept that attempts to unify the two architectures, it is worth pointing out that using the two models for the two categories of objects is able to help us generate goal poses for various free-floating and articulated objects. Moreover, by finetuning pretrained models from a weighted SVD combination, we are able to outperform the models on their respective training datasets categories. In future work, we wish to generalize the mathematics of the combined architecture beyond tasks that involve articulated and free-floating objects. We would also like to explore how such a unified architecture can aid motion planning as a geometric suggestor.